Please help me with this assignment.- Physics, Engineering Science.


  1. a. A motorcyclist traveling with a constant speed of 15m/s passes a school crossing corner, where the speed limit is 10m/s. Just as the motorist passes the school crossing sign, a police officer on a motorcycle stopped there, starts in pursuit with a constant acceleration of 3.0m/s2. I.    How much time elapses before the officer meets the motorist? 

II. What is the Police officer’s speed at that time? 

III. At that time, what distance has each vehicle traveled? 

 

b. The volume of an object as a function of time is calculated by 

V= At3+ B/t

, where t is time measured in seconds and V is in cubic meters. Determine the dimension of the constants A and B.


1
Expert's answer
2021-04-01T01:07:10-0400

(I) a. Let's assume that police officer catches the biker after t seconds

Distance travelled by them in time t is same

"15×t=u+\frac{1}{2}a×t^2"

Initial speed of police officer(u)=0 m/s, acceleration (a)=3m/s2

"15×t=\frac{1}{2}×3×t^2"

Time(t)=10 seconds


(II) The officer speed (v)=a×t=3×10=30m/s


(III) Distance (d)=0.5×a×t2=0.5×3×100=150 meters


b. Unit of (A×t3)= Unit of ("\frac{B}{t})" =Unit of (Volume)

Therefore,

Unit of (A)=(m3)/(sec3)

Unit of (B)=(m3×sec)